# AT2018fyk: Candidate Tidal Disruption Event by a (Super)massive Black Hole Binary

### v1

@@ 1 Introduction
When a star passes too close to a supermassive black hole (SMBH), the strong gravitational field can cause the star to be torn apart in what is known as a tidal disruption event (TDE; Hills, 1975; Rees, 1988). A significant fraction of the resulting debris is then accreted onto the SMBH, causing a bright flare of radiation that can last for months to years. The mass fallback of the debris typically follows a power law, decaying at late times as $t^{-5/3}$ (Phinney, 1989). The energy dissipated by the fallback debris powers a luminous flare that in principle enables the measurement of the SMBH mass and even spin(Mockler et al., 2019; Ryu et al., 2020; Wen et al., 2020; Mummery et al., 2023b), although existing efforts at parameter estimation all involve significant assumptions. For this reason, TDEs have evolved into a powerful tool for investigating the properties of SMBHs.
@@
A stellar disruption by the SMBHB secondary is also possible and would be the exclusive channel for a TDE if the primary SMBH mass exceeds $10^{8}M_{\odot}$, which would swallow instead of disrupt stars (Hills, 1975). Although the probability of disruption by the secondary is low (proportional to the mass ratio) compared to that of the primary (Coughlin et al., 2017), the event rate associated with the secondary is shown to be significantly increased by the eccentric Kozai-Lidov (EKL) mechanism (Kozai, 1962; Lidov, 1962; Naoz, 2016a), as reported by Mockler et al. (2023). Such events may be more distinguishable observationally compared to cases of disruption by the primary. They not only exhibit enhanced variability in the light curve due to more pronounced perturbations from the primary, but may also produce huge discrepancies in SMBH mass estimates, specifically between estimates (i) based on emission from the TDE accretion disk, and (ii) those from host galaxy scaling relationships.
@@ 2.1 X-ray Spectra
Table 1: A log of the Swift and XMM-Newton observations of AT2018fyk used in this paper. The number of source counts in the photon 0.3–10 keV energy band is given. The spectra are background subtracted. For XMM epochs, the exposed time and counts are shown in order of cameras pn, MOS1, and MOS2, respectively.
@@
For the XMM-Newton data, we run the default SAS v20 (20211130) tools under the HEASOFT ftools software version 6.30 to extract source spectra and to filter the data. The EPIC pn detector (Strüder et al., 2001; Turner et al., 2001) provides the highest count rates but we also used the MOS1 and MOS2 data where available. We filtered the pn data for periods of enhanced background radiation by deselecting episodes where the 10–12 keV detection rate of pattern 0 events is higher than 0.4 counts s-1. We filtered the MOS1 & MOS2 data for periods of enhanced background radiation by deselecting episodes where the detection rate of pattern 0 events at energies $>10$ keV is higher than 0.35 counts s-1. The effective exposure time for each observation after filtering is given in Table 1. The observations are done with the pn and MOS1&2 detectors in Large Window (XMM 1) or Full Frame (XMM 2 & 4) mode, respectively. We do not use the XMM 3 observation obtained on April 13-2020 (MJD 58952.20), as the source was undetected (Wevers et al., 2021).
@@ 2.4 Hubble Space Telescope Imaging
The location of AT2018fyk was imaged by the Hubble Space Telescope on 19 August 2021 as part of programme 16239 (PI Foley). At this epoch observations were obtained in F225W and F275W with exposure times of 780 and 710 seconds, respectively. Since only two dither positions were obtained in each filter there remain numerous cosmic-ray hits which cannot fully be removed. However, a source at the location of AT2018fyk is clearly detected in both drizzled images obtained from the MAST archive (given the small number of dithers, a re-processing of the data to drizzle to a smaller pixel scale is not plausible). We measure the source magnitude in these epochs in a small (0.12 arcsecond) aperture centred on the position, obtaining magnitudes of F225W=24.48 $\pm$ 0.04 and F275W = 22.70 $\pm$ 0.02, when corrected for encircled energy. The use of a larger (0.32 arcsec) aperture gives magnitudes approximate 0.4 and 0.2 magnitudes brighter (again after correction for encircled energy) indicating there may be some, extended underlying emission.
@@ 3.1 X-ray spectral fitting
Throughout this paper, we fit the spectra using XSPEC version 12.12.0 (Arnaud, 1996) applying Poisson statistics (Cash 1979; C-stat in XSPEC). We quote parameter errors at a $1\sigma$ $(68.3\%)$ confidence level (CL), using the method $\rm{Statistic}=\rm{Statistic_{best-fit}}+\Delta\,Cstat$ (Arnaud, 1996) and assuming $\Delta{\rm Cstat}=1.0$ and $\Delta{\rm Cstat}=2.3$ for single and two parameter models, respectively (Wen et al., 2021). We use the slim disk model, slimd (Wen et al., 2022), to model the disk spectrum and thermally comptonized continuum model, thcomp (Zdziarski et al., 2020), to model the hard part of the spectrum. The priors of their parameters are listed in Table 2. We set a prior on the inclination as $[2^{\circ},60^{\circ}]$, where $0^{\circ}$ represents a face-on orientation. When the disk is nearly edge-on ($>60^{\circ}$), a low accretion rate disk spectrum can be well fitted by a high accretion rate disk model, due to self-shielding by the finite aspect ratio of the disk (Wen et al., 2020). Edge-on TDEs typically exhibit additional absorption due to the sight line crossing outflowing material or intersecting the thick disc (Dai et al., 2018; Wen et al., 2020). However, there is no evidence for the presence of additional absorption in addition to the (Galactic) absorption parameterized by $N_{\rm H}=1.15\times 10^{20}\mathrm{cm^{-2}}$ (Wevers et al., 2021). Therefore, we enforce an upper limit on the disk inclination of $60^{\circ}$. We provide more details about the slimd model and its capability to constrain disk parameters in Appendix B.
@@ 3.1.1 Swift 1 Spectrum: Inner Accretion Disk
Wevers et al. (2021) shown that the source starts in a thermally dominated X-ray spectral state, which later develops to a power-law dominated state $\sim$100–200 days after the UV/optical peak. When analysing the time-averaged Swift spectra, we found that the power-law component becomes increasingly important after the date 2018/10/17. As a result, we designate the time-averaged spectrum from 2018/09/22 to 2018/10/17 as Swift 1. We fit the spectrum with a slim disk model (slimd) (Wen et al., 2022), a thin disk model (diskbb), and a simplified but general disk model valid in the Wien emission limit (tdediscspec) (Mummery, 2021; Mummery et al., 2023b), each attenuated by an absorption model TBAbs with a Galactic column density of NH = 1.15×$\times 10^{20}$ cm-2 (the same as in Wevers et al. 2021). Table 3 lists the fit results. See Figure 1 for the spectrum with the best-fit slim disk model overlaid. For the slim disk model (slimd), the free parameters are the BH mass ($M_{\bullet}$), the dimensionless BH spin ($a_{\bullet}$), the disk inclination ($\theta$), and the disk accretion rate $\dot{m}$ (in units of Eddington accretion rate, $\dot{M}_{\rm Edd}=1.37\times 10^{21}M_{6}~{}{\rm kg~{}s}^{-1}$, where $M_{6}=M_{\bullet}/(10^{6}M_{\odot})$). We find that the spectrum can be well fitted with a $\rm{Cstat/d.o.f}\approx 1.0$ for all the three models. These results show that the early X-ray spectrum originated from the inner part of an accretion disk. The inferred values for the disk bolometric luminosity and $M_{\bullet}$, as determined through the slim disk model fits, are $1.4^{+4.0}_{-0.8}\times 10^{43}$ erg/s and $0.5^{+1.5}_{-0.3}\times 10^{5}M_{\odot}$ at $1\sigma$ CL, respectively. The disk bolometric luminosity is consistent with the Eddington luminosity of $L_{\rm Edd}\approx 2.5\times 10^{43}$ erg/s for a BH mass of $M_{\bullet}=2\times 10^{5}M_{\odot}$, indicating that the disk is in a super-Eddington phase (also supported by the high accretion rate of Swift 1 in Table 3). The best-fit unabsorbed X-ray luminosity is $L_{\rm X}=7.5\times 10^{41}$ erg/s (for the band 0.3–10 keV), which implies that the ratio $L_{\rm X}/L_{\rm d}$ is in the range of $\approx 1-13\%$.
Table 3: The best-fit parameter values and their $1\sigma$ errors from X-ray spectral fits to the Swift 1 time-averaged spectrum. We fit the spectrum with either the fit-function TBAbs $\times$ slimd , TBAbs $\times$ diskbb, or TBAbs $\times$ tdediscspec. For the TBAbs model, we fix the NH to the value 1.15$\times$ 1020 cm-2. Note that the Eddington accretion rate is defined as $\dot{M}_{\rm Edd}=1.37\times 10^{21}M_{6}~{}{\rm kg~{}s}^{-1}$.
@@ 3.1.2 Evolution of Power-Law Component
We fit the Swift and XMM-Newton X-ray spectra simultaneously, with the fit-function (TBAbs $\times$ Thcomp(slimd)) consisting of a slim disk model and a thermally comptonized continuum (Zdziarski et al., 2020), absorbed by a Galactic column density of NH = 1.15$\times$ 1020 cm-2. For the slim disk component (slimd), we force $M_{\bullet}$, $a_{\bullet}$ and $\theta$ to have time-independent values, but we allow for different accretion rates $\dot{m}$ at each epoch. We do allow the (time-independent) $M_{\bullet}$, $a_{\bullet}$ and $\theta$ parameters to float freely and hence the values can be different from the values determined from the fit to the Swift 1 and XMM 2 spectra separately. We will discuss this in more detail in Section 3.1.3 and Appendix A.
We first fit the spectra by allowing the different epochs to have a different optical depth and a different electron temperature for the Thcomp model. We get a good fit with the total ${\rm Cstat/d.o.f.}=6749.3/6713$. We then fit the spectra by requiring all the epochs to have the same electron temperature. We also get a good fit with the total ${\rm Cstat/d.o.f.}=6755.7/6722$. The unfolded data and best-fit model spectrum for each epoch are shown in Figure 1. The best-fit parameters are listed in Table 4. The $\Delta{\rm AIC}=-11.6$ 222 $\Delta{\rm AIC}=\Delta Cstat+2\Delta K$, where $K$ is the number of free parameters. Generally speaking, when comparing the best-fit statistics of two models, $\Delta{\rm AIC}=5$ and $10$ are considered to be strong and very strong evidence, respectively, against the weaker model. (Akaike, 1974), obtained when comparing the fit where the electron temperatures are allowed to be different with the fit where they are tied together, shows that using the same corona electron temperature for all the epochs is strongly preferred by the data. We can see from Figure 1 that the contribution of the hard component to the overall spectrum starts to become larger about one month after the discovery on 2018 September 8 (MJD = 58369.2), a trend which continues for the next three months. After epoch 4, the parameters of the best-fit spectral model stay roughly constant. This can for instance be seen from the fitted optical depth of each epoch. The optical depth grows from 0.1 at Swift 2 to about 0.3 at Swift 4. We interpreted these results as the formation of a corona starting about one month after the peak of the UV emission. Its optical depth increases until it reaches a constant value in about 100 days.
@@ 3.1.3 Fitted Black Hole Mass and Spin
We also estimate $M_{\bullet}$ with the Swift-1 fits using a thin disk and a model for a spreading TDE disk (Mummery et al., 2023b). Assuming the disk terminates at the innermost stable circular orbit (ISCO) radius $r_{\rm ISCO}$, the thin disk model predicts the evolution of the maximum disk temperature as $T_{\rm eff,max}\approx 78(\frac{r_{\rm ISCO}}{r_{g}})^{-3/4}(\frac{\dot{m}}{M_{7}\eta_{-1}})^{1/4}~{}\rm{eV}$ (Shakura & Sunyaev, 1973), where $r_{g}=GM_{\bullet}/c^{2}$, $M_{7}=M_{\bullet}/10^{7}M_{\odot}$, $\eta_{-1}=\eta/0.1$, $G$ is Newton’s constant, $c$ is the speed of light and $\eta$ is the radiative efficiency). We get the mean effective temperature $T_{\rm eff,mean}\approx 2\sqrt{\frac{r_{\rm ISCO}}{r_{\rm out}}}T_{\rm eff,max}$ (for the case of $r_{\rm out}\gg r_{\rm ISCO}$) by integrating the flux across radii. Due to electron scattering in a hot atmosphere with finite temperature gradients, the real X-ray flux will be higher than the corresponding black-body flux (Shimura & Takahara, 1995), resulting in a higher effective temperature (by a factor of $f_{c}$, the spectral hardening factor).
@@ 3.2.1 Swift UVOT and HST observations
Past studies show that TDE UV/optical light curves (after peak) decay initially as a power-law (typically $<350$ days), then reaching a plateau. The initial power-law decay is often consistent with theoretical fallback rates, although the precise early-time emission mechanism remains unclear (Loeb & Ulmer, 1997; Piran et al., 2015; Roth et al., 2020). Late-time emission in the plateau phase is thought to originate from the outer edge of a circularized accretion disk (Ulmer, 1999; van Velzen et al., 2019; Hammerstein et al., 2023; Mummery et al., 2023a). However, in AT2018fyk, both the decay during the first 500 days and the decay after the rebrightening are well-fit by power-laws, with no evidence of a plateau. The outer disk UV emission is thus probably fainter than the observed values. Using the slim disk model (Wen et al., 2023), we can calculate the expected $w1$ magnitude from the outer disk using the parameters derived from fitting the X-ray spectra. For a disk around a BH with $M_{\bullet}=2\times 10^{5}M_{\odot}$, and an accretion rate of $\dot{m}=2$, we get $m_{w1}\sim 24$, which is indeed much fainter than the observed value of $2016<m_{w1}<$ 20 and also far below the host brightness ($m_{w1}\sim 21$) (Wevers et al., 2019). As the brightness of the disk in $w1$ is fainter than that of the host, we predict that the observed $w1$ emission will continue to decay following a power-law until it reaches the host level, which is expected to occur approximately 5000 days after discovery. However, if the star was disrupted by the secondary of a SMBHB, accretion to the primary BH would eventually interrupt the power-law decay.

### v3

@@ 1 Introduction
When a star passes too close to a supermassive black hole (SMBH), the strong gravitational field can cause the star to be torn apart in what is known as a tidal disruption event (TDE; Hills, 1975; Rees, 1988). A significant fraction of the resulting debris is then accreted onto the SMBH, causing a bright flare of radiation that can last for months to years. The mass fallback of the debris typically follows a power law, decaying at late times as $t^{-5/3}$ (Phinney, 1989). The energy dissipated by the fallback debris powers a luminous flare that in principle enables the measurement of the SMBH mass and even spin(Mockler et al., 2019; Ryu et al., 2020; Wen et al., 2020; Mummery et al., 2023), although existing efforts at parameter estimation all involve significant assumptions. For this reason, TDEs have evolved into a powerful tool for investigating the properties of SMBHs.
@@
A stellar disruption by the SMBHB secondary is also possible and would be the exclusive channel for a TDE if the primary SMBH mass exceeds $10^{8}M_{\odot}$, which would swallow instead of disrupt stars (Hills, 1975). Although the probability of disruption by the secondary is low (proportional to the mass ratio) compared to that of the primary (Coughlin et al., 2017), the event rate associated with the secondary is shown to be significantly increased by the eccentric Kozai-Lidov (EKL) mechanism (Kozai, 1962; Lidov, 1962; Naoz, 2016), as reported by Mockler et al. (2023). Such events may be more distinguishable observationally compared to cases of disruption by the primary. They not only exhibit enhanced variability in the light curve due to more pronounced perturbations from the primary, but may also produce huge discrepancies in SMBH mass estimates, specifically between estimates (i) based on emission from the TDE accretion disk, and (ii) those from host galaxy scaling relationships.
@@ 2.1 X-ray Spectra
Table 1: A log of the Swift and XMM-Newton observations of AT2018fyk used in this paper. The number of source counts in the photon 0.3–10 keV energy band is given. The spectra are background subtracted. For XMM epochs, the exposure time and counts are shown in order of cameras pn, MOS1, and MOS2, respectively.
@@
For the XMM-Newton data, we run the default SAS v20 (20211130) tools under the HEASOFT ftools software version 6.30 to extract source spectra and to filter the data. The EPIC pn detector (Strüder et al., 2001; Turner et al., 2001) provides the highest count rates but we also use the MOS1 and MOS2 data where available. We filter the pn data for periods of enhanced background radiation by deselecting episodes where the 10–12 keV detection rate of pattern 0 events is higher than 0.4 counts s-1. We filter the MOS1 & MOS2 data for periods of enhanced background radiation by deselecting episodes where the detection rate of pattern 0 events at energies $>10$ keV is higher than 0.35 counts s-1. The effective exposure time for each observation after filtering is given in Table 1. The observations are done with the pn and MOS1&2 detectors in Large Window (XMM 1) or Full Frame (XMM 2 & 4) mode, respectively. We do not use the XMM 3 observation obtained on April 13, 2020 (MJD 58952.20), as the source was undetected (Wevers et al., 2021).
@@ 2.4 Hubble Space Telescope Imaging
The location of AT2018fyk was imaged by the Hubble Space Telescope on August 19, 2021 as part of programme 16239 (PI Foley). At this epoch, observations were obtained in F225W and F275W with exposure times of 780 and 710 seconds, respectively. Since only two dither positions were obtained in each filter, there remain numerous cosmic-ray hits which cannot be fully removed. However, a source at the location of AT2018fyk is clearly detected in both drizzled images obtained from the MAST archive (given the small number of dithers, a re-processing of the data to drizzle to a smaller pixel scale is not plausible). We measure the source magnitude in these epochs in a small (0.12 arcsecond) aperture centred on the position, obtaining magnitudes of F225W=24.48 $\pm$ 0.04 and F275W = 22.70 $\pm$ 0.02, when corrected for encircled energy. The use of a larger (0.32 arcsec) aperture gives magnitudes approximately 0.4 and 0.2 magnitudes brighter (again after correction for encircled energy), indicating there may be some extended underlying emission.
@@ 3.1 X-ray spectral fitting
Throughout this paper, we fit the spectra using XSPEC version 12.12.0 (Arnaud, 1996), applying Poisson statistics (Cash 1979; C-stat in XSPEC). We quote parameter errors at a $1\sigma$ $(68.3\%)$ confidence level (CL), using the method $\rm{Statistic}=\rm{Statistic_{best-fit}}+\Delta\,Cstat$ (Arnaud, 1996) and assuming $\Delta{\rm Cstat}=1.0$ and $\Delta{\rm Cstat}=2.3$ for single and two parameter models, respectively (Wen et al., 2021). We use the slim disk model, slimd (Wen et al., 2022), to model the disk spectrum and the thermally comptonized continuum model, thcomp (Zdziarski et al., 2020), to model the hard part of the spectrum. The priors of their parameters are listed in Table 2. We set a prior on the inclination as $[2^{\circ},60^{\circ}]$, where $0^{\circ}$ represents a face-on orientation. When the disk is nearly edge-on ($>60^{\circ}$), a low accretion rate disk spectrum can be well fitted by a high accretion rate disk model, due to self-shielding by the finite aspect ratio of the disk (Wen et al., 2020). Edge-on TDEs typically exhibit additional absorption due to the sight line crossing outflowing material or intersecting the thick disk (Dai et al., 2018; Wen et al., 2020). However, there is no evidence for the presence of additional absorption in addition to the (Galactic) absorption parameterized by $N_{\rm H}=1.15\times 10^{20}\mathrm{cm^{-2}}$ (Wevers et al., 2021). Therefore, we enforce an upper limit on the disk inclination of $60^{\circ}$. We provide more details about the slimd model and its capability to constrain disk parameters in Appendix B.
@@ 3.1.1 Swift 1 Spectrum: Inner Accretion Disk
Wevers et al. (2021) showed that the source starts in a thermally dominated X-ray spectral state, which later develops into a power-law dominated state $\sim$100–200 days after the UV/optical peak. When analysing the time-averaged Swift spectra, we found that the power-law component becomes increasingly important after the date 2018/10/17. As a result, we designate the time-averaged spectrum from 2018/09/22 to 2018/10/17 as Swift 1. We fit the spectrum with a slim disk model (slimd) (Wen et al., 2022), a thin disk model (diskbb), and a simplified but general disk model valid in the Wien emission limit (tdediscspec) (Mummery, 2021; Mummery et al., 2023), each attenuated by an absorption model TBAbs with a Galactic column density of NH = 1.15×$\times 10^{20}$ cm-2 (the same as in Wevers et al. 2021). Table 3 lists the fit results. See Figure 1 for the spectrum with the best-fit slim disk model overlaid. For the slim disk model (slimd), the free parameters are the BH mass ($M_{\bullet}$), the dimensionless BH spin ($a_{\bullet}$), the disk inclination ($\theta$), and the disk accretion rate $\dot{m}$ (in units of Eddington accretion rate, $\dot{M}_{\rm Edd}=1.37\times 10^{21}M_{6}~{}{\rm kg~{}s}^{-1}$, where $M_{6}=M_{\bullet}/(10^{6}M_{\odot})$). We find that the spectrum can be well fitted with a $\rm{Cstat/d.o.f}\approx 1.0$ for all the three models. These results show that the early X-ray spectrum originated from the inner part of an accretion disk. The inferred values for the disk bolometric luminosity and $M_{\bullet}$, as determined through the slim disk model fits, are $1.4^{+4.0}_{-0.8}\times 10^{43}$ erg/s and $0.5^{+1.5}_{-0.3}\times 10^{5}M_{\odot}$ at $1\sigma$ CL, respectively. The disk bolometric luminosity is consistent with the Eddington luminosity of $L_{\rm Edd}\approx 2.5\times 10^{43}$ erg/s for a BH mass of $M_{\bullet}=2\times 10^{5}M_{\odot}$, indicating that the disk is in a super-Eddington phase (also supported by the high accretion rate of Swift 1 in Table 3). The best-fit unabsorbed X-ray luminosity is $L_{\rm X}=7.5\times 10^{41}$ erg/s (for the band 0.3–10 keV), which implies that the ratio $L_{\rm X}/L_{\rm d}$ is in the range of $\approx 1-13\%$.
Table 3: The best-fit parameter values and their $1\sigma$ errors from X-ray spectral fits to the Swift 1 time-averaged spectrum. We fit the spectrum with either the fit-function TBAbs $\times$ slimd , TBAbs $\times$ diskbb, or TBAbs $\times$ tdediscspec. For the TBAbs model, we fix the NH to the value 1.15$\times$ 1020 cm-2. Note that, the Eddington accretion rate is defined as $\dot{M}_{\rm Edd}=1.37\times 10^{21}M_{6}~{}{\rm kg~{}s}^{-1}$.
@@ 3.1.2 Evolution of Power-Law Component
We fit the Swift and XMM-Newton X-ray spectra simultaneously, with the fit-function (TBAbs $\times$ Thcomp(slimd)) consisting of a slim disk model and a thermally comptonized continuum (Zdziarski et al., 2020), absorbed by a Galactic column density of NH = 1.15$\times$ 1020 cm-2. For the slim disk component (slimd), we force $M_{\bullet}$, $a_{\bullet}$ and $\theta$ to have time-independent values, while allowing for different accretion rates $\dot{m}$ at each epoch. We do allow the (time-independent) $M_{\bullet}$, $a_{\bullet}$ and $\theta$ parameters to float freely and hence the values can be different from the values determined from the fit to the Swift 1 and XMM 2 spectra separately. We will discuss this in more detail in Section 3.1.3 and Appendix A.
We first fit the spectra by allowing the different epochs to have a different optical depth and a different electron temperature for the Thcomp model. We get a good fit with the total ${\rm Cstat/d.o.f.}=6749.3/6713$. We then fit the spectra by requiring all the epochs to have the same electron temperature. We also get a good fit with the total ${\rm Cstat/d.o.f.}=6755.7/6722$. The unfolded data and best-fit model spectrum for each epoch are shown in Figure 1. The best-fit parameters are listed in Table 4. The $\Delta{\rm AIC}=-11.6$ 222ΔAIC=$\Delta{\rm AIC}=\Delta Cstat+2\Delta K$, where $K$ is the number of free parameters. Generally speaking, when comparing the best-fit statistics of two models, $\Delta{\rm AIC}=5$ and $10$ are considered to be strong and very strong evidence, respectively, against the weaker model. (Akaike, 1974), obtained when comparing the fit where the electron temperatures are allowed to be different with the fit where they are tied together, shows that using the same corona electron temperature for all the epochs is strongly preferred by the data. We can see from Figure 1 that the contribution of the hard component to the overall spectrum starts to become larger about one month after the discovery on 2018 September 8 (MJD = 58369.2), a trend which continues for the next three months. After epoch 4, the parameters of the best-fit spectral model stay roughly constant. This can, for instance, be seen from the fitted optical depth of each epoch. The optical depth grows from 0.1 at Swift 2 to about 0.3 at Swift 4. We interpret these results as the formation of a corona starting about one month after the peak of the UV emission. Its optical depth increases until it reaches a constant value in about 100 days.
@@ 3.1.3 Fitted Black Hole Mass and Spin
We also estimate $M_{\bullet}$ with the Swift-1 fits using a thin disk and a model for a spreading TDE disk (Mummery et al., 2023). Assuming the disk terminates at the innermost stable circular orbit (ISCO) radius $r_{\rm ISCO}$, the thin disk model predicts the evolution of the maximum disk temperature as $T_{\rm eff,max}\approx 78(\frac{r_{\rm ISCO}}{r_{g}})^{-3/4}(\frac{\dot{m}}{M_{7}\eta_{-1}})^{1/4}~{}\rm{eV}$ (Shakura & Sunyaev, 1973), where $r_{g}=GM_{\bullet}/c^{2}$, $M_{7}=M_{\bullet}/10^{7}M_{\odot}$, $\eta_{-1}=\eta/0.1$, $G$ is Newton’s constant, $c$ is the speed of light, and $\eta$ is the radiative efficiency). We get the mean effective temperature $T_{\rm eff,mean}\approx 2\sqrt{\frac{r_{\rm ISCO}}{r_{\rm out}}}T_{\rm eff,max}$ (for the case of $r_{\rm out}\gg r_{\rm ISCO}$) by integrating the flux across radii. Due to electron scattering in a hot atmosphere with finite temperature gradients, the real X-ray flux will be higher than the corresponding black-body flux (Shimura & Takahara, 1995), resulting in a higher effective temperature (by a factor of $f_{c}$, the spectral hardening factor).
@@ 3.2.1 Swift UVOT and HST observations
Past studies show that TDE UV/optical light curves (after peak) decay initially as a power-law (typically $<350$ days), then reach a plateau. The initial power-law decay is often consistent with theoretical fallback rates, although the precise early-time emission mechanism remains unclear (Loeb & Ulmer, 1997; Piran et al., 2015; Roth et al., 2020). Late-time emission in the plateau phase is thought to originate from the outer edge of a circularized accretion disk (Ulmer, 1999; van Velzen et al., 2019; Hammerstein et al., 2023; Mummery et al., 2024). However, in AT2018fyk, both the decay during the first 500 days and the decay after the rebrightening are well-fit by power-laws, with no evidence of a plateau. The outer disk UV emission is thus probably fainter than the observed values. Using the slim disk model (Wen et al., 2023), we can calculate the expected $w1$ magnitude from the outer disk using the parameters derived from fitting the X-ray spectra. For a disk around a BH with $M_{\bullet}=2\times 10^{5}M_{\odot}$, and an accretion rate of $\dot{m}=2$, we get $m_{w1}\sim 24$, which is indeed much fainter than the observed value of $2016<m_{w1}<$ 20 and also far below the host brightness ($m_{w1}\sim 21$) (Wevers et al., 2019). As the brightness of the disk in $w1$ is fainter than that of the host, we predict that the observed $w1$ emission will continue to decay following a power-law until it reaches the host level, which is expected to occur approximately 5000 days after discovery. However, if the star was disrupted by the secondary of a SMBHB, accretion to the primary BH would eventually interrupt the power-law decay.